what is the conic section modeled by the equation 6x^2+14x=75-9y^2
Answer by lwsshak3(11628)

what is the conic section modeled by the equation:
6x^2+14x=75-9y^2
6x^2+14x+9y^2=75
complete the square
6(x^2+(7/3)x+49/36)+9y^2=75+49/6
6(x+7/6)^2+9y^2=499/6
This is an equation of an ellipse with horizontal major axis and center at (-7/6,0)